### v2

@@ 2 Methods
The MAMMAL method is built around three core components: the model architecture, the molecular prompt syntax, and extensive pretraining. In Subsection 2.1, we detail the architecture and its enhancements to the standard transformer framework. Subsection 2.2 focuses on the molecular prompt syntax, a key feature that enables the support of a diverse range of pretraining and downstream tasks for drug discovery. Finally, Subsection 2.3 outlines the pretraining process, which fascilitates leveraging large, cross-domain datasets and handling multiple entities simultaneously.
@@ 2.1 MAMMAL Architecture
The MAMMAL framework builds on the transformer architecture introduced by Vaswani et al. [19], and is inspired by the T5 framework [20] to formulate tasks as sequence-to-sequence problems within a unified model. MAMMAL introduces three primary features:
  • Task modeling in either an encoder-only mode, akin to BERT [21], or an encoder-decoder autoregressive mode [19]. The weights of the encoder stack are shared across both modes, enabling multi-task training that integrates tasks from both types. Model parameter updates are performed through gradient accumulation across all tasks.
  • Integral support for the molecular prompt syntax through the new Modular Tokenizer component, which facilitates the extension of molecular domain vocabularies and the incorporation of new domain vocabularies without necessitating the retraining of existing models.
  • Supporting numerical values (scalar) as both inputs being fed into the model, and also as outputs that the model can learn to predict. This is done in a continuous way, not requiring any binning or translation to a discrete space.
More details on the architecture and how scalars inputs and outputs are supported can be found in Appendix A
@@ 2.3 Pretraining
MAMMAL is designed as a comprehensive foundation model, capable of spanning multiple domains and accommodating a variety of entities. It is intended to support diverse task types, ranging from representation-focused tasks to generation-oriented ones. To achieve this, MAMMAL is trained on multiple tasks concurrently. Pretraining was conducted on 2 billion samples sourced from six datasets, which are all publicly available, covering three distinct domains across seven tasks. Table 1 summarizes these tasks, detailing the relevant domains, entity types, and specific datasets. Additional details about the pretraining are provided in Appendix C.
@@
”Number of Samples” describes the post filtering number of samples actually used. A single model
@@
was pretrained with all of the listed tasks, accumulating knowledge spanning multiple domains.
2.4 Evaluation
3 Results
Below, we present each benchmark used to evaluate ibm/biomed.omics.bl.sm.ma-ted-458m, along with performance results from the corresponding fine-tuned models. Each benchmark description includes background on the task, its significance for drug discovery, relevant prior models, and data statistics. A summary of the benchmarks, along with SOTA and MAMMAL results, is presented in Table 2 and visualized in Figure 1(E). Examples of encoder inputs and decoder labels for each benchmark are provided in Table LABEL:table_task_queries.
Table 2: Comparison of SOTA and MAMMAL Performance Across Benchmarks
Benchmark Domain Type Metric SOTA MAMMAL Imp.
Cell type GE cls $\uparrow$ F1 [28] 0.710 0.763±0.012 7.5 %
BBBP SM cls $\uparrow$ AUROC [29] 0.937 0.957±0.006 2.2 %
ClinTox SM cls $\uparrow$ AUROC [29] 0.948 0.986±0.007 4.0 %
DTI Prot.+SM reg $\downarrow$ NRMSE 0.942±0.028 [36] 0.906±0.011 3.8 %
00footnotetext: For NRMSE lower is better. For other metrics (AUROC, CDRH3-AAR, Pearson, Spearman, and F1) higher is better. Each row shows results from a MAMMAL model fine-tuned from ibm/biomed.omics.bl.sm.ma-ted-458m for the corresponding task. Abbreviations: in ”Type” column: ”cls” = classification, ”reg” = regression, ”gen” = generation. ”Imp.” = improvement (percentage) of our model over SOTA. In ”Domain” column: ”GE” = genes expression, ”SM” = small molecule, ”Prot.” = protein.
3.1 Cell Type Annotation
Cell type prediction enables researchers to distinguish between different cell populations, such as those associated with various diseases [9, 10, 11, 12]. It is also crucial for understanding how diseases or drugs affect different cell types. In recent years, a variety of methods have been developed for this task, including approaches based on marker genes, correlation-based techniques, and annotation using classification [37]. Recent advances in transformer-based and large-scale foundation models [38, 28, 39] have improved performance by utilizing the full list of genes as input, in contrast to earlier methods where gene selection was based on highly variable genes.
The input for this task commonly consists of gene expression (GE) values from single-cell RNA-seq data. The benchmark we used is based on the Zheng68k dataset [40], which is derived from human peripheral blood mononuclear cells and is widely used for evaluating cell-type annotation performance. The dataset contains 68,579 cells across 11 cell types and originally included 32,738 genes. Preprocessing involved normalization, log-transformation of expression values and followed filtering out non-expressed genes, leaving around 20,387 genes. Similar to the approach in [41], our model uses a ranked list of expressed gene names, ordered by their expression levels, as input. The label to predict is provided in the cell ontology format “CL:NNNNNN” (see Table LABEL:table_task_queries).
3.2 BBBP and ClinTox
Drugs must meet various criteria regarding both efficacy and safety. In this study, we selected two relevant benchmarks from MoleculeNet [42], a widely used suite of benchmarks for small-molecule drugs: BBBP and ClinTox. The task in the BBBP benchmark is to predict the ability of drugs to penetrate the blood-brain barrier, a crucial factor in the development of drugs targeting the central nervous system. The ClinTox benchmark involves two related tasks: (1) predicting failure in clinical toxicity trials, and (2) predicting FDA approval status. The overall performance on ClinTox is reported as the average performance across these two tasks.
MoLFormer [29], a well-established model for molecular embeddings trained on 1.1 billion SMILES sequences, has achieved state-of-the-art performance on both the BBBP and ClinTox benchmarks. In our study, we adopted the benchmarks from [29], which provided predefined splits for training, validation, and testing. As shown in  2, MAMMAL surpassed MoLFormer on both benchmarks, achieving an average AUROC of 0.937 on BBBP and 0.986 on ClinTox - representing improvements of 2.2% and 4%, respectively, over the state of the art.
3.3 Cancer-Drug Response (CDR)
For our study, we used three subsets of the GDSC database: GDSC1 and GDSC2, available through the Therapeutics Data Commons (TDC) [47], and referred in the paper as Cancer-Drug Response 1 and Cancer-Drug Response 2 respectively; and a subset published in [31], referred as Cancer-Drug Response 3. Table S3 summarizes the number of cell lines, drugs, and cell-drug pairs in these datasets. We used the random splits provided by TDC for Cancer-Drug Response 1 and 2, while for Cancer-Drug Response 3, we followed the split methodology outlined in [31], reserving 5% of the data for the test set, stratified by TCGA [48] pathways associated with the cancer cell lines.
During fine-tuning we used only gene-expression profiles and SMILES representations of drugs, as shown in the example query in Table LABEL:table_task_queries. Similar to the input format for cell type annotation, gene-expression profiles were provided as ranked lists of gene names based on their expression levels. For predicting continuous IC50 values, MAMMAL was utilized in regression mode, taking advantage of its built-in support for floating point scalar predictions. As demonstrated in Table 2, our model outperforms the current SOTA models for Cancer-Drug Response 1 and 2, achieving a 3.4% increase in Pearson correlation values. Additionally, it yields results comparable to the SOTA for the Cancer-Drug Response 3 benchmark, with a slight improvement of 0.5%..
3.4 Targeted Antibody Design
Antibodies are a family of proteins produced by the immune system to neutralize foreign antigens and are of particular interest due to their high specificity and strong binding to target molecules [49, 50]. These characteristics have made them a crucial class of therapeutics, driving significant research efforts into the design of new antibody-based drug candidates [51, 52, 53, 5]. Antigen-binding fragments (Fabs) of antibodies consist of two amino acid chains, referred to as the heavy and light chains. Each chain is further divided into four framework (FR) regions and three complementarity-determining regions (CDRs). While FR regions are typically conserved, CDRs exhibit significant variation in their amino acid composition and are generally the primary determinants of binding affinity to the target antigen. When designing novel antibodies for a specific antigen, the typical approach is to explore alternative CDRs that could produce a new, functional antibody with high binding affinity to the target [49, 50, 54, 32].
Recently, several deep learning methods have been developed for targeted antibody design, framing CDR prediction as an infilling task [54, 55, 56, 57, 58, 32, 59]. These models predict missing CDR regions, represented by MASK tokens, using the amino acid sequences of both the antigen and the antibody’s FR regions. While prior approaches often rely on structural data, this information is scarce and challenging to obtain [60]. In contrast, we fine-tune MAMMAL for the targeted antibody design task using only the sequence data from the antigen and the antibody’s FR regions.
3.5 Antibody-Antigen Binding Prediction
The HER2 dataset was divided into train (70%), validation (15%) and test (15%) sets. Fine-tuning involved three concurrent tasks: mask infilling for the antibody heavy chains, and two classification tasks for antibody-antigen binding prediction: one utilizing the heavy chain sequence and the other based on the CDR3 subsequence. We focus on the heavy chain classification task for performance evaluation. As depicted in table 2, our model achieved an average AUC of 0.879, slightly surpassing the SOTA, which incorporated structural data that our model did not.
3.6 T-Cell Receptor-Epitope Binding
[…]
3.7 Protein Protein Interaction - $\Delta$$\Delta$G Prediction
$\Delta G=kT\ln(K_{D}),$ (1)
where $k$ is the Boltzmann constant and $T$ is the temperature [63].
$\Delta\Delta G=\Delta G_{\text{mutant}}-\Delta G_{\text{wild-type}}.$
The SKEMPI dataset [63] provides information on changes in thermodynamic parameters, including $\Delta G$, and kinetic rate constants due to mutations in protein-protein (PP) complexes whose structures are available in the Protein Data Bank [67]. This dataset is extensively utilized in the literature for predicting the effects of mutations on binding affinity, particularly in the context of $\Delta\Delta G$. A subset of SKEMPI comprising 1,131 samples of single mutations, S1131, is widely used. We adopt S1131 as our benchmark for predicting protein-protein $\Delta\Delta G$ and follow the common practice of reporting 10-fold cross-validation performance on this subset. The input query for our model includes the reference “wild-type” version of the complex and the corresponding mutated version, comprising only sequence data without any structural information. We leverage MAMMAL’s support for floating-point scalars to predict continuous $\Delta\Delta G$ in a regression task setting. Performance results are presented in Table 2. As shown in Table 2, our model achieved an average Pearson correlation of 0.852, significantly exceeding the previous sequence-only SOTA of 0.663. Compared to models that incorporate structural data, our model’s performance remains competitive, falling just 1.6% short of the SOTA performance of 0.866 [64].
3.8 Drug-Target Interaction
4 Discussion
Artificial intelligence (AI) holds great promise for transforming the drug discovery process by enhancing efficiency, accuracy, and speed. Developed with this goal in mind, our proposed method enables the creation of AI models capable of handling diverse tasks across multiple data domains. MAMMAL reformulates tasks as sequence-to-sequence problems and introduces several key architectural enhancements: support for both encoder-only and encoder-decoder modes, a multi-domain extensible syntax for inputs and outputs, and the direct handling of numerical values through continuous token embeddings. MAMMAL has been applied to pretrain model ibm/biomed.omics.bl.sm.ma-ted-458m while aligning inputs across diverse datasets - including small molecule, protein, antibody, and gene expression data-using a variety of pretraining tasks. This multi-align approach enables the integration of cross-domain pharmaceutical knowledge into a single model and facilitates effective transfer learning to fine-tuned models for downstream applications. Demonstrated state-of-the-art performance of the multi-align fine-tuned models across diverse tasks, spanning multiple data domains and stages of the drug discovery pipeline, strongly supports the potential of the MAMMAL approach.
Supporting both encoder-only and encoder-decoder modes is motivated by the observation that different tasks benefit from distinct architectures: generative tasks with variable output lengths are well-suited to an encoder-decoder approach, while classification and regression tasks align more effectively with an encoder-only setup. Sharing the encoder stack across both modes optimizes it jointly across tasks, allowing each to benefit from the collective knowledge embedded within the encoder. Additionally, the extensible molecular prompt language enables flexible task formulation across diverse domains, unlike models constrained by fixed prompt structures. This flexibility enhances applicability across various domains, making it a versatile tool for researchers from different fields and promoting interdisciplinary collaboration.
Drug discovery is often hindered by a lack of large, high-quality datasets, particularly in biomedicine, where data generation and collection are costly and face challenges related to privacy, lack of standardization, and ethical constraints. A planned extension of our approach involves the support of free-text input, enabling pretraining on extensive biomedical text sources like PubMed as well as the incorporation of free-text segments into prompts of fine-tuned models. Free text provides rich, nuanced context and allows users greater flexibility beyond predefined formats. The integration of large language models, which capture human narratives, with our newly introduced multi-align model-demonstrated to effectively encode complex interactions within protein complexes, between proteins and small molecules, and between small molecules and cells, including their gene expression levels-holds promise of creating a model capable of uncovering new scientific knowledge.
We are pleased to announce the open-source release of the the code and the pretrained model weights for ibm/biomed.omics.bl.sm.ma-ted-458m, inviting the community to explore, apply, and contribute to its ongoing development. By making MAMMAL accessible, we aim to foster collaboration that enriches the platform through the addition of new pretraining tasks and the integration of new domains, such as DNA sequences. We believe that such collaborative efforts will strengthen and enhance the capabilities of MAMMAL, broadening its applicability in the field of biomedicine and ultimately leading to new discoveries in drug research.
@@ 5 Data Availability
Cancer-Drug Response 1 and 2. GDSC1 and GDSC2 benchmarks were accessed with random splits from the TDC library (https://pypi.org/project/PyTDC/). Cancer-Drug Response 3. Benchmark obtained from DeepCDR [31] git repository (https://github.com/kimmo1019/DeepCDR/tree/master/data).
DTI. Benchmark was published by [36] and available in https://torchdrug.ai/docs/api/datasets.html#bindingdb
Ab Infilling. Data is taken from [32], which provides a preprocessed subset of the publicly available SAbDab database [60]. The preprocessing pipeline includes a similarity-based clustering for the data splits and sample level filtering which excludes samples that are considered in valid in [32]. For additional information, we refer to [32] and the publicly available codebase, https://github.com/THUNLP-MT/dyMEAN.
@@
Pretraining. ibm/biomed.omics.bl.sm.ma-ted-458m was pre-trained over OAS [23], UniProt [71], Zinc [24], PubChem [25], STRING [27] and CELLxGENE [26]. Appendix C describes the pre-processing steps applied.
@@ 6 Code Availability
The model architecture, fine-tuning framework, and end-to-end examples are publicly available at https://github.com/BiomedSciAI/biomed-multi-alignment. This repository provides comprehensive resources for utilizing the model, including instructions for fine-tuning, and performing inference on various tasks. The pretrained model weights and tokenizer can be accessed via the Hugging Face model hub at https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m.
@@ Appendix A Architecture - additional details
Figure 2: A prompt, consisting of both token ids and scalars is processed and enters the encoder. Both the encoder and the decoder output logits which are used for classification loss. Additionally, the output of the encoder is sent to a (learned) scalars prediction head which allows to predict scalars for any subset of the tokens, and is used in the regression loss. In this illustration, a single scalar input (”12.7”) is being used, and a single scalars outputs are predicted by the model (”97.2”). However, the method fully supports an arbitrary number of input scalars and outputs.
One of the key aspects in MAMMAL method is the built in support for scalars inputs and outputs. Figure 2 illustrates how this is achieved.
A user prompt, usually expressed as a single text line, is processed into two input sequences: a. Input token IDs, which is a sequence of integer values representing tokens in the vocabulary, and b. a sequence of inputs scalars (by convention, containing NaNs for positions for which no input scalar is provided).
The input tokens ids are transformed using a learned token embedding, and the input scalars are transformed using a learned linear transformation which projects each single scalar element into the model dimension (e.g. 768).
@@
For scalars outputs (gene expression, binding free enery, etc.) the encoder stack has an additional prediction head, which outputs a scalar value for every input element. How to deal with scalars outputs in locations that there is no scalar label is up to the user choice, but the default is to ignore those.
The support of scalars outputs in the encoder-decoder mode is an improvement that we intend to add in future generations of the model/method.

### v3

@@ 3 Methods
The MAMMAL method is built around three core components: The model architecture, the molecular prompt syntax, and extensive pretraining. In Subsection 3.1, we detail the architecture and its enhancements to the standard transformer.
Subsections 3.2 and 3.3 focus on the molecular representations and prompt syntax, key features that enable the support of a diverse range of pretraining and downstream tasks for drug discovery. Finally, Subsection 3.4 outlines the pretraining process, which facilitates leveraging large, cross-domain datasets and handling multiple entities simultaneously.
@@ 3.1 MAMMAL Architecture
The MAMMAL architecture builds upon the transformer architecture introduced by Vaswani et al. [75] and draws inspiration from the T5 framework [76] while introducing several key modifications on top of it. This design conceptualizes tasks as sequence-to-sequence problems within a unified model, introducing three key features:
Optimized for Representation and Generation Tasks
Biomedical models for drug discovery must perform diverse tasks, ranging from token-level or prompt-level representation to tasks requiring strong generative capabilities. For instance, [77] introduced a non-causal, decoder-only protein language model optimized for both understanding and generating protein-related data, outperforming state-of-the-art baselines. MAMMAL is designed to jointly optimize for such diverse tasks, enabling the model to learn from multiple stages of the drug discovery pipeline. This collective learning is facilitated by a shared encoder, which bridges gaps between distinct AI applications in drug discovery. Specifically, MAMMAL supports both encoder-only and encoder-decoder autoregressive modes. Encoder-only mode excels in representation-heavy tasks, such as classification and regression, while encoder-decoder mode is better suited for generative tasks. By sharing encoder stack weights across these modes, MAMMAL facilitates efficient multi-task training, with parameter updates conducted through gradient accumulation across all tasks.
Flexible Multi-Domain Structured Prompts
Drug discovery tasks often require prompts that describe molecular complexes composed of multiple entities. These prompts may also include attributes related to the entire complex or specific internal entities to improve prediction accuracy. To address this, MAMMAL employs a modular tokenizer that enables multi-domain structured prompts by assigning distinct sub-tokenizers to each entity type. Compared to free-text prompts, structured prompts provide a consistent and explicit representation of input data, capturing relationships and attributes between entities, and facilitating the training. While MAMMAL primarily focuses on structured input, incorporating free text as an additional data source is a promising future direction. Recent advancements in biomedical large language models [78, 79, 27], demonstrate the potential of free-text prompts to enhance versatility and expand training datasets. Future iterations of MAMMAL could integrate free text directly, mix it with structured prompts, or leverage embeddings derived from extensively pre-trained text models [80, 81]. These embeddings could harness the latent knowledge encoded within external models to enrich the system’s performance.
Numerical Values Integration
A notable innovation of MAMMAL is its native support for numerical values (e.g., 2, 3.14, 10,000.1) as both inputs and outputs. Numerical data is crucial for modeling drug discovery tasks and allows for direct representation of important attributes. This capability expands the range of tasks MAMMAL can handle, such as binding affinity prediction and protein folding. Many biomedical large language models either lack support for numerical data [78], employ limited solutions such as binning [27], or rely on digit-based representations [79], which can inflate input length. [82] offer compelling digit-based approach of splitting the number into semantic parts (mainly digits), supporting arbitrary scalar values while adding only a limited number of tokens to the vocabulary. While this simplifies integration with standard language models, it inflates the number of input/output tokens significantly. This is especially evident in tasks like gene expression prediction which require supporting thousands of scalars in a single prompt and expected model output. [83] suggests a new loss term that considers numerical proximity, with the main advantage of not requiring any additional prediction head for scalars, allowing to introduce numerical values support within existing language model architectures. However, it is currently limited to a fixed vocabulary, not allowing support for arbitrary (possibly unseen in train time) scalar values. Other methods, such as discretization, demonstrated in [84] through pre-trained VQ-VAE [85] that encodes and decodes 3D spatial positions, represent another approach limited to a specific use case. MAMMAL, however, integrates continuous numerical values directly into its embedding space via a projection layer. The resulting embeddings align with the input token embeddings, ensuring simple, efficient, and effective utilization of numerical data.
Further details about the MAMMAL architecture can be found in Appendix A.
3.2 Entity Representation
Each entity domain supported by MAMMAL makes use of a representation approach that has been selected as follows:
  • Small Molecules. Represented using SMILES sequences (Simplified Molecular Input Line Entry System). For example, paracetamol is represented as CC(=O)NC1=CC=C(O)C=C1.
  • Gene Expression. Represented as an ordered list of gene names, sorted in descending order based on log-normalized and binned expression values. In case of ties, genes are sorted alphabetically. This representation applies to both single-cell and bulk RNA expression data.
  • Proteins. Represented as concatenated amino acid chains, preserving original source-data order. This representation contains only sequence data, without structural information.
  • Antibodies. Represented similarly to proteins, using concatenated amino acid sequences. Each chain is prefixed with a token indicating its type (heavy or light).
@@ 3.3 Prompt Syntax
The prompt syntax is built around a modular tokenizer architecture that governs how molecular data is interpreted and encoded. At its core, the tokenizer defines a common syntax using special tags that represent molecular entities, sequences, attributes, and interactions across diverse molecular systems. It supports multi-domain inputs by delegating different segments of the input sequence to specialized sub-tokenizers, each responsible for handling a specific data type or modality.
@@
The prompt syntax applies uniformly to both inputs and outputs of the model. It is also designed to be extensible: New tags can be introduced into individual sub-tokenizers or added to the shared token set without disrupting existing functionality. Thanks to this modular structure, updates to domain-specific sub-tokenizers or the token vocabulary maintain backward compatibility, enabling interoperability between newly trained and previously deployed models. More detailed explanations and examples of the prompt syntax can be found in Appendix B.
@@ 3.4 Pretraining
MAMMAL is designed as a comprehensive foundation model, capable of spanning multiple domains and accommodating a variety of entities. It is intended to support diverse task types, ranging from representation-focused tasks to generation-oriented ones. To achieve this, MAMMAL is trained on multiple tasks concurrently. Pretraining was conducted on two billion samples sourced from six datasets, which are all publicly available, covering three distinct domains across seven tasks. Table 2 summarizes these tasks, detailing the relevant domains, entity types, and specific datasets. Additional details about the pretraining are provided in Appendix C.
@@
”Number of Samples” lists the post-filtering number of samples actually used. A single model was pretrained with all of the listed tasks, accumulating knowledge spanning multiple domains.
@@ 4 Data Availability
Cancer-Drug Response 1 and 2. The GDSC1 and GDSC2 benchmarks were accessed with random splits from the TDC library (https://pypi.org/project/PyTDC/). Cancer-Drug Response 3. The benchmark was obtained from the DeepCDR [28] git repository (https://github.com/kimmo1019/DeepCDR/tree/master/data).
DTI. This benchmark was published by [33] and is available in https://torchdrug.ai/docs/api/datasets.html#bindingdb
Ab Infilling. This data was taken from [29], which provides a preprocessed subset of the publicly available SAbDab database [57]. The preprocessing pipeline includes a similarity-based clustering for the data splits and sample-level filtering that excludes samples considered invalid in [29]. For additional information, we refer to [29] and the publicly available codebase: https://github.com/THUNLP-MT/dyMEAN.
@@
Pretraining. The ibm/biomed.omics.bl.sm.ma-ted-458m was pre-trained over OAS [86], UniProt [93], Zinc [87], PubChem [88], STRING [90] and CELLxGENE [89]. Appendix C describes the pre-processing steps applied.
@@ 5 Code Availability
The model architecture, fine-tuning framework, and end-to-end examples are publicly available at: https://github.com/BiomedSciAI/biomed-multi-alignment. This GitHub repository provides comprehensive resources for working with the model, including setup instructions, fine-tuning guidelines, and inference workflows for multiple downstream tasks.
Pretrained model weights and the associated tokenizer are hosted on the Hugging Face Model Hub and can be accessed at: https://huggingface.co/ibm/biomed.omics.bl.sm.ma-ted-458m.
Additionally, selected fine-tuned model checkpoints and tokenizers are available via: https://huggingface.co/models?other=base_model:finetune:ibm-research/biomed.omics.bl.sm.ma-ted-458m, and can be explored interactively through the Hugging Face Space at: https://huggingface.co/spaces/ibm/biomed-multi-alignment.
6 Acknowledgments
Partially Supported by IBM-Technion Research Collaboration
@@ Appendix A Architecture - Additional Details
Figure 3: A prompt, consisting of both token IDs and scalars is processed and enters the encoder. Both the encoder and the decoder output logits which are used for classification loss. Additionally, the output of the encoder is sent to a (learned) scalars prediction head which allows the prediction scalars for any subset of the tokens, and is used in the regression loss. In this illustration, a single scalar input (”12.7”) is being used, and a single scalar’s outputs are predicted by the model (”97.2”). However, the method fully supports an arbitrary number of input scalars and outputs.
One of the key aspects in the MAMMAL method is the built-in support for scalar inputs and outputs. Figure 3 illustrates how this is achieved.
A user prompt, usually expressed as a single text line, is processed into two input sequences: a. Input token IDs, which are a sequence of integer values representing tokens in the vocabulary, and b. a sequence of inputs scalars (by convention, containing NaNs for positions for which no input scalar is provided).
The input tokens IDs are transformed using a learned token embedding, and the input scalars are transformed using a learned linear transformation that projects each single scalar element into the model dimension (e.g, 768).
@@
For scalars outputs (gene expression, binding free energy, etc.), the encoder stack has an additional prediction head, which outputs a scalar value for every input element. How to deal with scalar outputs in locations where there is no scalar label is the user’s choice, but the default is to ignore those.
The support of scalar outputs in the encoder-decoder mode is an improvement that we intend to add in future generations of the model/method.